OpenStudy (anonymous):

each exterior of a regular polygon measures 45 deggrees. find the sum of the measures of the polygons interior angles. part1. what is the sum of the measures of the exterior angles of any polygon

OpenStudy (anonymous):

360/45 = 8 sides Interior 180-45 = 135 135 * 8 = 1080 degrees. sum of the measures of the exterior angles of any polygon - 360 degrees